# Vision Mamba: Efficient Visual Representation Learning with Bidirectional State Space Model

### v2

@@ Abstract
Recently the state space models (SSMs) with efficient hardware-aware designs, i.e., the Mamba deep learning model, have shown great potential for long sequence modeling. Meanwhile building efficient and generic vision backbones purely upon SSMs is an appealing direction. However, representing visual data is challenging for SSMs due to the position-sensitivity of visual data and the requirement of global context for visual understanding. In this paper, we show that the reliance on self-attention for visual representation learning is not necessary and propose a new generic vision backbone with bidirectional Mamba blocks (Vim), which marks the image sequences with position embeddings and compresses the visual representation with bidirectional state space models. On ImageNet classification, COCO object detection, and ADE20k semantic segmentation tasks, Vim achieves higher performance compared to well-established vision transformers like DeiT, while also demonstrating significantly improved computation & memory efficiency. For example, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory when performing batch inference to extract features on images with a resolution of 1248$\times$1248. The results demonstrate that Vim is capable of overcoming the computation & memory constraints on performing Transformer-style understanding for high-resolution images and it has great potential to be the next-generation backbone for vision foundation models.
Figure 1: Performance and efficiency comparisons between DeiT [59] and our Vim model. For the accuracy comparison, we first pretrain DeiT and Vim on IN1K classification dataset [9], then we finetune the generic backbones on different downstream dense prediction tasks, i.e., semantic segmentation, object detection, instance segmentation. Results show that the proposed Vim outperforms DeiT on both pretraining and finetuning tasks. Vim is also more computation and memory efficient than DeiT in dealing with high-resolution images. For example, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory when performing batch inference to extract features on images with a resolution of 1248$\times$1248, i.e., 6084 tokens per image.
[…]
@@ 1 Introduction
Recent research advancements have led to a surge of interest in the state space model (SSM). Originating from the classic Kalman filter model [29], modern SSMs excel at capturing long-range dependencies and benefit from parallel training. Some SSM-based methods, such as the linear state-space layers (LSSL) [21], structured state space sequence model (S4) [20], diagonal state space (DSS) [23], and S4D [22], are proposed to process sequence data across a wide range of tasks and modalities, particularly on modeling long-range dependencies. They are efficient in processing long sequences because of convolutional computation and near-linear computation. 2-D SSM [2], SGConvNeXt [36], and ConvSSM [51] combine SSM with CNN or Transformer architecture to process 2-D data. The recent work, Mamba [19], incorporates time-varying parameters into the SSM and proposes a hardware-aware algorithm to enable very efficient training and inference. The superior scaling performance of Mamba indicates that it is a promising alternative to Transformer in language modeling. Nevertheless, a generic pure-SSM-based backbone network has not been explored for processing visual data, such as images and videos.
Vision Transformers (ViTs) have achieved great success in visual representation learning, excelling in large-scale self-supervised pre-training and high performance on downstream tasks. Compared with convolutional neural networks, the core advantage lies in that ViT can provide each image patch with data/patch-dependent global context through self-attention. This differs from convolutional networks that use the same parameters, i.e., the convolutional filters, for all positions. Another advantage is the modality-agnostic modeling by treating an image as a sequence of patches without 2D inductive bias, which makes it the preferred architecture for multimodal applications [3, 35, 39]. At the same time, the self-attention mechanism in Transformers poses challenges in terms of speed and memory usage when dealing with long-range visual dependencies, e.g., processing high-resolution images.
@@
Compared with other SSM-based models for vision tasks, Vim is a pure-SSM-based method and models images in a sequence manner, which is more promising for a generic and efficient backbone. Thanks to the bidirectional compressing modeling with positional awareness, Vim is the first pure-SSM-based model to handle dense prediction tasks. Compared with the most convincing Transformer-based model, i.e., DeiT [59], Vim achieves superior performance on ImageNet classification. Furthermore, Vim is more efficient in terms of GPU memory and inference time for high-resolution images. The efficiency in terms of memory and speed empowers Vim to directly perform sequential visual representation learning without relying on 2D priors (such as the 2D local window in ViTDet [37]) for high-resolution visual understanding tasks while achieving higher accuracy than DeiT.
@@ 2 Related Work
Architectures for generic vision backbone. In the early eras, ConvNet [33] serves as the de-facto standard network design for computer vision. Many convolutional neural architectures [32, 55, 50, 24, 56, 62, 25, 71, 57, 49] have been proposed as the vision backbone for various visual applications. The pioneering work, Vision Transformer (ViT) [13] changes the landscape. It treats an image as a sequence of flattened 2D patches and directly applies a pure Transformer architecture. The surprising results of ViT on image classification and its scaling ability encourage a lot of follow-up works [60, 58, 61, 15]. One line of works focuses on hybrid architecture designs by introducing 2D convolutional priors into ViT [68, 8, 14, 12]. PVT [65] proposes a pyramid structure Transformer. Swin Transformer [41] applies self-attention within shift windows. Another line of works focuses on improving traditional 2D ConvNets with more advanced settings [66, 40]. ConvNeXt [42] reviews the design space and proposes pure ConvNets, which can be scalable as ViT and its variants. RepLKNet [11] proposes to scale up the kernel size of existing ConvNets to bring improvements.
@@
Though these dominant follow-up works demonstrate superior performance and better efficiency on ImageNet [9] and various downstream tasks [38, 73] by introducing 2D priors, with the surge of large-scale visual pretraining [1, 16, 5] and multi-modality applications [48, 34, 35, 39, 3, 28], vanilla Transformer-style model strikes back to the center stage of computer vision. The advantages of larger modeling capacity, unified multi-modality representation, being friendly to self-supervised learning etc., make it the preferred architecture. However, the number of visual tokens is limited due to the quadratic complexity of Transformer. There are plenty of works [7, 64, 31, 6, 10, 47, 54] to address this long-standing and prominent challenge, but few of them focus on visual applications. Recently, LongViT [67] built an efficient Transformer architecture for computational pathology applications via dilated attention. The linear computation complexity of LongViT allows it to encode the extremely long visual sequence. In this work, we draw inspiration from Mamba [19] and explore building a pure-SSM-based model as a generic vision backbone without using attention, while preserving the sequential, modality-agnostic modeling merit of ViT.
@@
State space models for long sequence modeling. [20] proposes a Structured State-Space Sequence (S4) model, a novel alternative to CNNs or Transformers, to model the long-range dependency. The promising property of linearly scaling in sequence length attracts further explorations. [52] proposes a new S5 layer by introducing MIMO SSM and efficient parallel scan into S4 layer. [17] designs a new SSM layer, H3, that nearly fills the performance gap between SSMs and Transformer attention in language modeling. [45] builds the Gated State Space layer on S4 by introducing more gating units to improve the expressivity. Recently, [19] proposes a data-dependent SSM layer and builds a generic language model backbone, Mamba, which outperforms Transformers at various sizes on large-scale real data and enjoys linear scaling in sequence length. In this work, we explore transferring the success of Mamba to vision, i.e., building a generic vision backbone purely upon SSM without attention.
@@
State space models for visual applications. [26] uses 1D S4 to handle the long-range temporal dependencies for video classification. [46] further extends 1D S4 to handle multi-dimensional data including 2D images and 3D videos. [27] combines the strengths of S4 and self-attention to build TranS4mer model, achieving state-of-the-art performance for movie scene detection. [63] introduces a novel selectivity mechanism to S4, largely improving the performance of S4 on long-form video understanding with a much lower memory footprint. [72] supplants attention mechanisms with a more scalable SSM-based backbone to generate high-resolution images and process fine-grained representation under affordable computation. [44] proposes U-Mamba, a hybrid CNN-SSM architecture, to handle the long-range dependencies in biomedical image segmentation. The above works either apply SSM to specific visual applications or build a hybrid architecture by combining SSM with convolution or attention. Different from them, we build a pure-SSM-based model, which can be adopted as a generic vision backbone.
@@ 3 Method
The goal of Vision Mamba (Vim) is to introduce the advanced state space model (SSM), i.e., Mamba [19], to computer vision. This section begins with a description of the preliminaries of SSM. It is followed by an overview of Vim. We then detail how the Vim block processes input token sequences and proceed to illustrate the architecture details of Vim. The section concludes with an analysis of the efficiency of the proposed Vim.
@@ 3.1 Preliminaries
The SSM-based models, i.e., structured state space sequence models (S4) and Mamba are inspired by the continuous system, which maps a 1-D function or sequence $x(t)\in\mathbb{R}\mapsto y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{\mathtt{N}}$. This system uses $\mathbf{A}\in\mathbb{R}^{\mathtt{N}\times\mathtt{N}}$ as the evolution parameter and $\mathbf{B}\in\mathbb{R}^{\mathtt{N}\times 1}$, $\mathbf{C}\in\mathbb{R}^{1\times\mathtt{N}}$ as the projection parameters.
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (1)
$y(t)$ $=\mathbf{C}h(t).$
@@
After the discretization of $\mathbf{\overline{A}}$, $\mathbf{\overline{B}}$, the discretized version of Eq. (1) using a step size $\mathbf{\Delta}$ can be rewritten as:
@@ 3.2 Vision Mamba
where $\mathbf{t}_{p}^{\mathtt{j}}$ is the $\mathtt{j}$-th patch of $\mathbf{t}$, $\mathbf{W}\in\mathbb{R}^{(\mathtt{P}^{2}\cdot\mathtt{C})\times\mathtt{D}}$ is the learnable projection matrix. Inspired by ViT [13] and BERT [30], we also use class token to represent the whole patch sequence, which is denoted as $\mathbf{t}_{cls}$. We then send the token sequence ($\mathbf{T}_{\mathtt{l}-1}$) to the $\mathtt{l}$-th layer of the Vim encoder, and get the output $\mathbf{T}_{\mathtt{l}}$. Finally, we normalize the output class token $\mathbf{T}_{\mathtt{L}}^{0}$ and feed it to the multi-layer perceptron (MLP) head to get the final prediction $\hat{p}$, as follows:
$\mathbf{T}_{l}$ $=\mathbf{Vim{}}(\mathbf{T}_{\mathtt{l}-1})+\mathbf{T}_{\mathtt{l}-1},$ (6)
$\mathbf{f}$ $=\mathbf{Norm}(\mathbf{T}_{\mathtt{L}}^{0}),$
$\hat{p}$ $=\mathbf{MLP}(\mathbf{f}),$
where $\mathbf{Vim{}}$ is the proposed vision mamba block, $\mathtt{L}$ is the number of layers, and $\mathbf{Norm}$ is the normalization layer.

### v3

@@ Abstract
Recently the state space models (SSMs) with efficient hardware-aware designs, i.e., the Mamba deep learning model, have shown great potential for long sequence modeling. Meanwhile building efficient and generic vision backbones purely upon SSMs is an appealing direction. However, representing visual data is challenging for SSMs due to the position-sensitivity of visual data and the requirement of global context for visual understanding. In this paper, we show that the reliance on self-attention for visual representation learning is not necessary and propose a new generic vision backbone with bidirectional Mamba blocks (Vim), which marks the image sequences with position embeddings and compresses the visual representation with bidirectional state space models. On ImageNet classification, COCO object detection, and ADE20k semantic segmentation tasks, Vim achieves higher performance compared to well-established vision transformers like DeiT, while also demonstrating significantly improved computation & memory efficiency. For example, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory when performing batch inference to extract features on images with a resolution of 1248$\times$1248. The results demonstrate that Vim is capable of overcoming the computation & memory constraints on performing Transformer-style understanding for high-resolution images and it has great potential to be the next-generation backbone for vision foundation models. Code and models are released at https://github.com/hustvl/Vim
Figure 1: Performance and efficiency comparisons between DeiT (Touvron et al., 2021a) and our Vim model. Results show that Vim outperforms DeiT on both ImageNet classification and downstream detection and segmentation tasks and is more computation and memory efficient than DeiT in dealing with high-resolution images. For example, Vim is 2.8$\times$ faster than DeiT and saves 86.8% GPU memory when performing batch inference to extract features on images with a resolution of 1248$\times$1248, i.e., 6084 tokens per image.
@@ 1 Introduction
Recent research advancements have led to a surge of interest in the state space model (SSM). Originating from the classic Kalman filter model (Kalman, 1960), modern SSMs excel at capturing long-range dependencies and benefit from parallel training. Some SSM-based methods, such as the linear state-space layers (LSSL) (Gu et al., 2021b), structured state space sequence model (S4) (Gu et al., 2021a), diagonal state space (DSS) (Gupta et al., 2022), and S4D (Gu et al., 2022), are proposed to process sequence data across a wide range of tasks and modalities, particularly on modeling long-range dependencies. They are efficient in processing long sequences because of convolutional computation and near-linear computation. 2-D SSM (Baron et al., 2023), SGConvNeXt (Li et al., 2022b), and ConvSSM (Smith et al., 2023a) combine SSM with CNN or Transformer architecture to process 2-D data. The recent work, Mamba (Gu & Dao, 2023), incorporates time-varying parameters into the SSM and proposes a hardware-aware algorithm to enable very efficient training and inference. The superior scaling performance of Mamba indicates that it is a promising alternative to Transformer in language modeling. Nevertheless, a generic pure-SSM-based backbone network has not been explored for processing visual data, such as images and videos.
Vision Transformers (ViTs) have achieved great success in visual representation learning, excelling in large-scale self-supervised pre-training and high performance on downstream tasks. Compared with convolutional neural networks, the core advantage lies in that ViT can provide each image patch with data/patch-dependent global context through self-attention. This differs from convolutional networks that use the same parameters, i.e., the convolutional filters, for all positions. Another advantage is the modality-agnostic modeling by treating an image as a sequence of patches without 2D inductive bias, which makes it the preferred architecture for multimodal applications (Bavishi et al., 2023; Li et al., 2023; Liu et al., 2023). At the same time, the self-attention mechanism in Transformers poses challenges in terms of speed and memory usage when dealing with long-range visual dependencies, e.g., processing high-resolution images.
@@
Compared with other SSM-based models for vision tasks, Vim is a pure-SSM-based method and models images in a sequence manner, which is more promising for a generic and efficient backbone. Thanks to the bidirectional compressing modeling with positional awareness, Vim is the first pure-SSM-based model to handle dense prediction tasks. Compared with the most convincing Transformer-based model, i.e., DeiT (Touvron et al., 2021a), Vim achieves superior performance on ImageNet classification. Furthermore, Vim is more efficient in terms of GPU memory and inference time for high-resolution images. The efficiency in terms of memory and speed empowers Vim to directly perform sequential visual representation learning without relying on 2D priors (such as the 2D local window in ViTDet (Li et al., 2022c)) for high-resolution visual understanding tasks while achieving higher accuracy than DeiT.
@@ 2 Related Work
Architectures for generic vision backbone. In the early eras, ConvNet (LeCun et al., 1998) serves as the de-facto standard network design for computer vision. Many convolutional neural architectures (Krizhevsky et al., 2012; Szegedy et al., 2015; Simonyan & Zisserman, 2014; He et al., 2016; Tan & Le, 2019; Wang et al., 2020a; Huang et al., 2017; Xie et al., 2017; Tan & Le, 2021; Radosavovic et al., 2020) have been proposed as the vision backbone for various visual applications. The pioneering work, Vision Transformer (ViT) (Dosovitskiy et al., 2020) changes the landscape. It treats an image as a sequence of flattened 2D patches and directly applies a pure Transformer architecture. The surprising results of ViT on image classification and its scaling ability encourage a lot of follow-up works (Touvron et al., 2021b; Tolstikhin et al., 2021; Touvron et al., 2022; Fang et al., 2022). One line of works focuses on hybrid architecture designs by introducing 2D convolutional priors into ViT (Wu et al., 2021; Dai et al., 2021; d’Ascoli et al., 2021; Dong et al., 2022). PVT (Wang et al., 2021) proposes a pyramid structure Transformer. Swin Transformer (Liu et al., 2021) applies self-attention within shift windows. Another line of works focuses on improving traditional 2D ConvNets with more advanced settings (Wang et al., 2023b; Liu et al., 2022a). ConvNeXt (Liu et al., 2022b) reviews the design space and proposes pure ConvNets, which can be scalable as ViT and its variants. RepLKNet (Ding et al., 2022) proposes to scale up the kernel size of existing ConvNets to bring improvements.
@@
Though these dominant follow-up works demonstrate superior performance and better efficiency on ImageNet (Deng et al., 2009) and various downstream tasks (Lin et al., 2014; Zhou et al., 2019) by introducing 2D priors, with the surge of large-scale visual pretraining (Bao et al., 2022; Fang et al., 2023; Caron et al., 2021) and multi-modality applications (Radford et al., 2021; Li et al., 2022a, 2023; Liu et al., 2023; Bavishi et al., 2023; Jia et al., 2021), vanilla Transformer-style model strikes back to the center stage of computer vision. The advantages of larger modeling capacity, unified multi-modality representation, being friendly to self-supervised learning etc., make it the preferred architecture. However, the number of visual tokens is limited due to the quadratic complexity of Transformer. There are plenty of works (Choromanski et al., 2021; Wang et al., 2020b; Kitaev et al., 2020; Child et al., 2019; Ding et al., 2023; Qin et al., 2023; Sun et al., 2023) to address this long-standing and prominent challenge, but few of them focus on visual applications. Recently, LongViT (Wang et al., 2023c) built an efficient Transformer architecture for computational pathology applications via dilated attention. The linear computation complexity of LongViT allows it to encode the extremely long visual sequence. In this work, we draw inspiration from Mamba (Gu & Dao, 2023) and explore building a pure-SSM-based model as a generic vision backbone without using attention, while preserving the sequential, modality-agnostic modeling merit of ViT.
@@
State space models for long sequence modeling. (Gu et al., 2021a) proposes a Structured State-Space Sequence (S4) model, a novel alternative to CNNs or Transformers, to model the long-range dependency. The promising property of linearly scaling in sequence length attracts further explorations. (Wang et al., 2022) proposes Bidirectional Gated SSM to replicate BERT (Devlin et al., 2018) results without attention. (Smith et al., 2023b) proposes a new S5 layer by introducing MIMO SSM and efficient parallel scan into S4 layer. (Fu et al., 2023) designs a new SSM layer, H3, that nearly fills the performance gap between SSMs and Transformer attention in language modeling. (Mehta et al., 2023) builds the Gated State Space layer on S4 by introducing more gating units to improve the expressivity. Recently, (Gu & Dao, 2023) proposes a data-dependent SSM layer and builds a generic language model backbone, Mamba, which outperforms Transformers at various sizes on large-scale real data and enjoys linear scaling in sequence length. In this work, we explore transferring the success of Mamba to vision, i.e., building a generic vision backbone purely upon SSM without attention.
@@
State space models for visual applications. (Islam & Bertasius, 2022) uses 1D S4 to handle the long-range temporal dependencies for video classification. (Nguyen et al., 2022) further extends 1D S4 to handle multi-dimensional data including 2D images and 3D videos. (Islam et al., 2023) combines the strengths of S4 and self-attention to build TranS4mer model, achieving state-of-the-art performance for movie scene detection. (Wang et al., 2023a) introduces a novel selectivity mechanism to S4, largely improving the performance of S4 on long-form video understanding with a much lower memory footprint. (Yan et al., 2023) supplants attention mechanisms with a more scalable SSM-based backbone to generate high-resolution images and process fine-grained representation under affordable computation. (Ma et al., 2024) proposes U-Mamba, a hybrid CNN-SSM architecture, to handle the long-range dependencies in biomedical image segmentation. The above works (Xing et al., 2024; Ma et al., 2024; Yan et al., 2023; Wang et al., 2023a; Islam et al., 2023; Nguyen et al., 2022; Islam & Bertasius, 2022) either apply SSM to specific visual applications or build a hybrid architecture by combining SSM with convolution or attention. Different from them, we build a pure-SSM-based model, which can be adopted as a generic vision backbone. It is noteworthy that VMamba (Liu et al., 2024), a concurrent work with our method, has demonstrated impressive results in visual recognition by incorporating Mamba with multi-directional scanning and a hierarchical network architecture. In contrast, Vim primarily concentrates on visual sequence learning and boasts a unified representation for multi-modality data.
@@ 3 Method
The goal of Vision Mamba (Vim) is to introduce the advanced state space model (SSM), i.e., Mamba (Gu & Dao, 2023), to computer vision. This section begins with a description of the preliminaries of SSM. It is followed by an overview of Vim. We then detail how the Vim block processes input token sequences and proceed to illustrate the architecture details of Vim. The section concludes with an analysis of the efficiency of the proposed Vim.
@@ 3.1 Preliminaries
The SSM-based models, i.e., structured state space sequence models (S4) and Mamba are inspired by the continuous system, which maps a 1-D function or sequence $x(t)\in\mathbb{R}\mapsto y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{\mathtt{N}}$. This system uses $\mathbf{A}\in\mathbb{R}^{\mathtt{N}\times\mathtt{N}}$ as the evolution parameter and $\mathbf{B}\in\mathbb{R}^{\mathtt{N}\times 1}$, $\mathbf{C}\in\mathbb{R}^{1\times\mathtt{N}}$ as the projection parameters. The continuous system works as follows: $h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t)$ and $y(t)=\mathbf{C}h(t)$.
@@
After the discretization of $\mathbf{\overline{A}}$, $\mathbf{\overline{B}}$, the discretized version using a step size $\mathbf{\Delta}$ can be rewritten as:
@@ 3.2 Vision Mamba
where $\mathbf{t}_{p}^{\mathtt{j}}$ is the $\mathtt{j}$-th patch of $\mathbf{t}$, $\mathbf{W}\in\mathbb{R}^{(\mathtt{P}^{2}\cdot\mathtt{C})\times\mathtt{D}}$ is the learnable projection matrix. Inspired by ViT (Dosovitskiy et al., 2020) and BERT (Kenton & Toutanova, 2019), we also use class token to represent the whole patch sequence, which is denoted as $\mathbf{t}_{cls}$. We then send the token sequence ($\mathbf{T}_{\mathtt{l}-1}$) to the $\mathtt{l}$-th layer of the Vim encoder, and get the output $\mathbf{T}_{\mathtt{l}}$. Finally, we normalize the output class token $\mathbf{T}_{\mathtt{L}}^{0}$ and feed it to the multi-layer perceptron (MLP) head to get the final prediction $\hat{p}$, as follows: $\mathbf{T}_{l}=\mathbf{Vim{}}(\mathbf{T}_{\mathtt{l}-1})+\mathbf{T}_{\mathtt{l}-1}$, $\mathbf{f}=\mathbf{Norm}(\mathbf{T}_{\mathtt{L}}^{0})$, and $\hat{p}=\mathbf{MLP}(\mathbf{f})$, where $\mathbf{Vim{}}$ is the proposed vision mamba block, $\mathtt{L}$ is the number of layers, and $\mathbf{Norm}$ is the normalization layer.
@@ 3.3 Vim Block
Specifically, we present the operations of Vim block in Algo. 29. The input token sequence $\mathbf{T}_{\mathtt{l}-1}$ is first normalized by the normalization layer. Next, we linearly project the normalized sequence to the $\mathbf{x}$ and $\mathbf{z}$ with dimension size $E$. Then, we process the $\mathbf{x}$ from the forward and backward directions. For each direction, we first apply the 1-D convolution to the $\mathbf{x}$ and get the $\mathbf{x}^{\prime}_{o}$. We then linearly project the $\mathbf{x}^{\prime}_{o}$ to the $\mathbf{B}_{o}$, $\mathbf{C}_{o}$, $\mathbf{\Delta}_{o}$, respectively. The $\mathbf{\Delta}_{o}$ is then used to transform the $\overline{\mathbf{A}}_{o}$, $\overline{\mathbf{B}}_{o}$, respectively. Finally, we compute the $\mathbf{y}_{forward}$ and $\mathbf{y}_{backward}$ through the SSM. The $\mathbf{y}_{forward}$ and $\mathbf{y}_{backward}$ are then gated by the $\mathbf{z}$ and added together to get the output token sequence $\mathbf{T}_{\mathtt{l}}$.
3.4 Architecture Details
In summary, the hyper-parameters of our architecture are listed as follows: $\mathtt{L}$ denotes the number of blocks, $\mathtt{D}$ denotes the hidden state dimension, $\mathtt{E}$ denotes the expanded state dimension, and $\mathtt{N}$ denotes the SSM dimension. Following ViT (Dosovitskiy et al., 2020) and DeiT (Touvron et al., 2021b), we first employ 16$\times$16 kernel size projection layer to get a 1-D sequence of non-overlapping patch embeddings. Subsequently, we directly stack $\mathtt{L}$ Vim blocks. By default, we set the number of blocks $\mathtt{L}$ to 24, SSM dimension $\mathtt{N}$ to 16. To align with the model sizes of DeiT series, we set the hidden state dimension $\mathtt{D}$ to 192 and expanded state dimension $\mathtt{E}$ to 384 for the tiny-size variant. For the small-size variant, we set $\mathtt{D}$ to 384 and $\mathtt{E}$ to 768.